OpenStudy (anonymous):

hepl me simplify sqaure root of 8 & square root of 7

OpenStudy (anonymous):

you mean \[\sqrt{8}\] and \[\sqrt{7}\]?

OpenStudy (anonymous):

you can do nothing with \[\sqrt{7}\]

OpenStudy (anonymous):

yeesss,

OpenStudy (anonymous):

as for \[\sqrt{8}\]you can say \[4\times 2=8\] so \[\sqrt{8}=\sqrt{4\times 2}=\sqrt{4}\times \sqrt{2}=2\sqrt{2}\]

OpenStudy (anonymous):

you see this fairly often so might be a good idea to remember that \[\sqrt{8}=2\sqrt{2}\] and also \[\sqrt{12}=2\sqrt{3}\]

OpenStudy (anonymous):

yess it is thanks

OpenStudy (anonymous):

yw